### v2

# Ithaca, NY 14853, USA

Abstract.
$m$-location is a local combinatorial condition for flag simplicial complexes introduced by Osajda. Osajda showed that simply connected $8$-located locally $5$-large complexes are hyperbolic. We treat the nonpositive curvature case of $7$-located locally $5$-large complexes.
We show that any minimal area disc diagram in a $7$-located locally $5$-large complex is itself $7$-located and locally $5$-large. We define a natural $\operatorname{CAT}(0)$ metric for $7$-located disc diagrams and use this to prove that simply connected $7$-located locally $5$-large complexes have quadratic isoperimetric function. Along the way, we prove that locally weakly systolic complexes are $7$-located locally $5$-large.
Key words and phrases:
7-located complex, combinatorial nonpositive curvature, weakly systolic complex
2010 Mathematics Subject Classification:
20F65, 20F67
1. Introduction
Metric and combinatorial notions of negative and nonpositive curvature have seen significant study and application in recent decades, especially in the areas of geometric group theory and metric graph theory [7, 5, 8, 10, 2, 4, 1, 9]. These notions are often given as local conditions that have global implications on spaces. Of particular note on the combinatorial side are the developments of $\operatorname{CAT}(0)$-cubical complexes (also known as median graphs) [7, 5], systolic complexes (also known as bridged graphs) [5, 8, 10] and Helly graphs [4].
The combinatorial nonpositive curvature condition of $m$-location for simplicial complexes was introduced by Osajda in [14] as a method for proving hyperbolicity for $3$-manifolds. While local systolicity of a complex depends only on the combinatorics around a single vertex, $m$-location depends on the combinatorics around pairs of adjacent vertices, allowing more flexibility. Osajda proved that simply connected $8$-located locally $5$-large complexes are Gromov hyperbolic and used this to obtain a new solution to a problem of Thurston. The second named author studied a version of $8$-location, suggested by Osajda [14, Subsection 5.1] and showed that simply connected, $8$-located simplicial complexes are Gromov hyperbolic [11]. In later work, the second named author introduced another combinatorial curvature condition, called the $5/9$-condition, and showed that the complexes which satisfy it, are $8$-located and therefore Gromov hyperbolic as well [12].
The current paper continues the study of $m$-located complexes. As noted by Osajda [14], systolic complexes are $7$-located. We strengthen this statement by showing that weakly systolic complexes are $7$-located. This provides many more examples of $7$-located locally $5$-large complexes, including thickenings of $\operatorname{CAT}(0)$ cube complexes whose links do not contain induced $4$-cycles [13].
Theorem A (Theorem 3.1).
Locally weakly systolic complexes are $7$-located locally $5$-large.
We also prove the following two theorems which show that $7$-location and local $5$-largeness have significant global consequences: disc diagrams with particularly nice combinatorial and geometric properties.
Theorem B (Theorem 5.9).
Minimal area disc diagrams in $7$-located locally $5$-large complexes are $7$-located locally $5$-large.
Let $P_{n}$ be the regular Euclidean $n$-gon subdivided into $n$ congruent triangles meeting at the barycenter of $P_{n}$. Let $T_{n}$ be a triangle of $P_{n}$.
Theorem C (Theorem 4.6).
Let $D$ be a $7$-located simplicial disc. There is a natural way to metrize the triangles of $D$ using only isometry types of $T_{4}$, $T_{5}$ and $T_{6}$ such that the resulting piecewise Euclidean triangle complex is $\operatorname{CAT}(0)$. The choice of metrization for each triangle depends only on the local combinatorics of $D$ around the triangle.
From Theorem B and Theorem C we obtain the following global coarse geometric and algebraic consequences.
Corollary D (Corollary 4.7).
A simply connected $7$-located locally $5$-large complex has quadratic isoperimetric function. Consequently, the fundamental group of a $7$-located locally $5$-large complex has quadratic isoperimetric function.
1.1. Structure of the paper
In Section 2 we present basic definitions and notation, in Section 3 we prove Theorem A, in Section 4 we prove Theorem C and in Section 5 we prove Theorem B.
2. Preliminaries
Let $X$ be a simplicial complex. If $v_{1},v_{2},\ldots,v_{k}$ are vertices of $X$ then we use the notation $\langle v_{1},v_{2},\ldots,v_{k}\rangle$ to denote the simplex spanned on the vertices $v_{1},v_{2},\ldots,v_{k}$, should it exist. A subcomplex $L$ of $X$ is full if any simplex of $X$ whose vertices are contained in $L$ is contained in $L$. For distinct vertices $v$ and $v^{\prime}$ of $X$ we write $v\sim v^{\prime}$ if $v$ and $v^{\prime}$ are adjacent in the $1$-skeleton of $X$ and otherwise we write $v\nsim v^{\prime}$. The simplicial complex $X$ is flag if any finite set of pairwise adjacent vertices of $X$ spans a simplex of $X$. Going forward we will always assume that $X$ is a flag complex.
Unless otherwise stated, when referring to distances between vertices of a simplicial complex we will mean the graph metric, i.e., the number of edges in the shortest $1$-skeleton path joining the vertices.
A cycle or loop $\gamma$ in $X$ is a subcomplex of $X$ isomorphic to a subdivision of $S^{1}$. We may denote a cycle $\gamma$ by a sequence $(v_{1},v_{2},\ldots,v_{k})$ where $\{v_{1},v_{2},\ldots,v_{k}\}$ is the set of vertices of $\gamma$ and where $\langle v_{i},v_{i+1}\rangle$ is an edge of $\gamma$ for each $i$, indices taken modulo $k$. The length of $\gamma$, denoted by $|\gamma|$, is the number of edges of $\gamma$. A $k$-wheel in $X$ is a subcomplex $W=(v_{0};v_{1},v_{2},\ldots,v_{k})$ of $X$ where $(v_{1},\ldots,v_{k})$ is a cycle and $\langle v_{0},v_{i},v_{i+1}\rangle$ is a triangle of $W$ for each $i$, indices taken modulo $k$. That is, the $k$-wheel $W$ is a simplicial cone on its boundary cycle $\partial W=(v_{1},v_{2},\ldots,v_{k})$.
Let $\gamma$ be a cycle in $X$. A disc diagram $(D,f)$ for $\gamma$ is a nondegenerate simplicial map $f:D\rightarrow X$ where $D$ is a triangulated disc, and $f|_{\partial D}$ maps $\partial D$ isomorphically onto $\gamma$. The disc diagram $(D,f)$ has minimal area if $D$ has the least possible number of triangles among all disc diagrams for $\gamma$.
By van Kampen’s Lemma, every nullhomotopic cycle in $X$ has a disc diagram. It is not difficult to see that when $X$ is a flag simplicial complex, any minimal area disc diagram in $X$ is also a flag simplicial complex.
Figure 1. A planar $(6,5)$-dwheel and a nonplanar $(6,5)$-dwheel.
A $(k,l)$-dwheel $W=W_{1}\cup W_{2}$ in $X$ is the union of two wheels $W_{1}=(w_{l};v_{1},v_{2},v_{3},\ldots,$ $v_{k})$ and $W_{2}=(v_{2};w_{1},w_{2},\ldots,w_{l})$ with $v_{3}=w_{l-1}$ and either $v_{1}=w_{1}$ or $v_{1}\sim w_{1}$. See Figure 1. If $v_{1}=w_{1}$, we call $W$ a planar dwheel. If $v_{1}\sim w_{1}$, we call $W$ a nonplanar dwheel. The boundary $\partial W$ of $W$ is the cycle $(w_{1},w_{2},\ldots,w_{l-1}=v_{3},v_{4},\ldots,v_{k})$. The boundary $\partial W$ is the subcomplex of $W$ induced on its vertices excluding the wheel centers $w_{l}$ and $v_{2}$. Note that $\partial W$ has length $k+l-4$ if $W$ is planar and $k+l-3$ if it is nonplanar.
The link of $X$ at a vertex $v$, denoted $X_{v}$, is the subcomplex of $X$ consisting of all simplices of $X$ that are disjoint from $v$ and that, together with $v$, span a simplex of $X$.
Definition 2.1.
A simplicial complex $X$ is $m$-located for $m\geq 4$ if it is flag and whenever a dwheel subcomplex $W=W_{1}\cup W_{2}$ of $X$ satisfies the conditions
  1. (1)
    $\partial W$ has length at most $m$, and
  2. (2)
    the wheels $W_{1}$ and $W_{2}$ are full subcomplexes of $X$,
the dwheel subcomplex $W$ is contained in the link $X_{v}$ of some vertex $v$.
Remark 2.2.
In the definition of $m$-locatedness, we may weaken the requirement that $W$ is contained in the link $X_{v}$ of a vertex by requiring only that $W$ is contained in the $1$-ball $B_{1}(v)$ centered at a vertex. This results in an equivalent definition since if $W$ is contained in $B_{1}(v)$ then either it is contained in $X_{v}$ or $v\in W$, which implies that the wheels of $W$ are not both full.
Definition 2.3.
Let $\sigma$ be a simplex of $X$. A flag simplicial complex is $k$-large if there are no full $j$-cycles in $X$, when $14\leq j\leq k-$ 1. We say $X$ is locally $k$-large if all its links are $k$-large. Note that if $X$ is $k$-large then it is locally $k$-large.
$a$$c$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$
Figure 2. An extended $5$-wheel.
For the purposes of this paper, we need only refer to local weak systolicity, a condition which together with simple connectedness is equivalent to weak systolicity by the local-to-global theorem of Chepoi and Osajda [6, Theorem A]. In order to define local weak systolicity, we first need to introduce some terminology. An extended $5$-wheel $\widehat{W}_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5};a)$ in a flag simplicial complex $X$ is a subcomplex consisting of a $5$-wheel $W_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5})$ together with an additional vertex $a$ that spans a triangle $\langle a,v_{1},v_{2}\rangle$ with $v_{1}$ and $v_{2}$ but that is not adjacent to any other vertex of $W_{5}$. See Figure 2.
Definition 2.4.
A flag simplicial complex $X$ is locally weakly systolic if it is $5$-large and every extended $5$-wheel is contained in the link of a vertex. This latter condition is called the $\widehat{W}_{5}$-condition.
3. Weakly systolic complexes are $7$-located
In this section we prove that locally weakly systolic complexes are $7$-located. Because locally weakly systolic complexes are $5$-large, they thus provide a large class of examples of $7$-located locally $5$-large complexes.
$w_{5}$$v_{1}=w_{1}$$v_{2}$$v_{3}=w_{4}$$v_{4}$$v_{5}$$w_{2}$$w_{3}$
Figure 3. A planar $(5,5)$-dwheel.
Theorem 3.1.
Let $X$ be a locally weakly systolic complex. Then $X$ is $7$-located.
Proof.
Let $W=W_{1}\cup W_{2}$ be a $(k,l)$-dwheel subcomplex of $X$ of boundary length at most $7$ such that $W_{1}$ and $W_{2}$ are full subcomplexes. By $5$-largeness of $X$, we have $k,l\geq 5$. Then, since $W$ has boundary length at most $7$, either $W$ is a planar or nonplanar $(5,5)$-dwheel or $W$ is a planar $(5,6)$-dwheel. By Remark 2.2 it suffices to show that $W$ is contained in the $1$-ball centered at a vertex of $X$.
Claim 3.2.
If $W=W_{1}\cup W_{2}$ is a planar $(5,5)$-dwheel then $W$ is contained in a $1$-ball of $X$.
Proof.
Let $W_{1}\cup W_{2}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5})\cup(v_{2};w_{1}=v_{1},w_{2},w_{3},w_{4}=v_{3},w_{5})$. Then $W_{1}$ and $W_{2}$ are full. See Figure 3.
Consider $W_{1}\cup\{w_{2}\}$. Note that $w_{2}\sim v_{1}$ and $w_{2}\sim v_{2}$. Also, since $W_{2}$ is a full $k$-wheel, the vertices $w_{2}$ and $w_{5}$ are distinct and $w_{2}\nsim w_{5}$ and $w_{2}\nsim v_{3}$. If we had $w_{2}\sim v_{4}$ then $(w_{2},v_{2},w_{5},v_{4})$ would be a $4$-cycle and then $5$-largeness of $X$ would contradict fullness of either $W_{1}$ or $W_{2}$. So $w_{2}\nsim v_{4}$. We also have $w_{2}\nsim v_{5}$ arguing otherwise similarly using the $4$-cycle $(w_{2},v_{2},w_{5},v_{5})$. Thus $W_{1}\cup\{w_{2}\}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5};w_{2})$ is an extended $5$-wheel. Thus, by the $\widehat{W_{5}}$-condition, there is a vertex $\bar{w}_{2}$ that is adjacent to but distinct from every vertex of $W_{1}\cup\{w_{2}\}$. If $\bar{w}_{2}=w_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\notin W_{1}\cup W_{2}$. By a similar argument, there is a vertex $\bar{w}_{3}\notin W_{1}\cup W_{2}$ that is adjacent to every vertex of $W_{1}\cup\{w_{3}\}$. If $\bar{w}_{2}=\bar{w}_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\neq\bar{w}_{3}$. Then $(v_{1},\bar{w}_{2},v_{3},\bar{w}_{3})$ is a $4$-cycle and $5$-largeness of $X$ and fullness of $W_{1}$ imply that $\bar{w}_{2}\sim\bar{w}_{3}$. Finally, the $5$-largeness of $X$ in consideration of the $4$-cycle $(\bar{w}_{2},w_{2},w_{3},\bar{w}_{3})$ implies that $W_{1}\cup W_{2}$ is contained in the $1$-ball centered at either $\bar{w}_{2}$ or $\bar{w}_{3}$. ∎
Claim 3.3.
If $W=W_{1}\cup W_{2}$ is a nonplanar $(5,5)$-dwheel then $W$ is contained in a $1$-ball of $X$.
Proof.
Let $W_{1}\cup W_{2}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5})$ $\cup(v_{2};w_{1},w_{2},w_{3},w_{4}=v_{3},w_{5})$. See Figure 4.
$w_{5}$$v_{1}$$v_{2}$$v_{3}=w_{4}$$v_{4}$$v_{5}$$w_{1}$$w_{2}$$w_{3}$
Figure 4. A nonplanar $(5,5)$-dwheel from the proof of Claim 3.3. The encircled vertices form an extended $5$-wheel $\widehat{W}_{1}$ contained in the link $X_{p}$ of the vertex $p$. The ensquared vertices form an extended $5$-wheel $\widehat{W}_{2}$ contained in $X_{q}$.
Let $\widehat{W}_{1}=W_{1}\cup\{w_{3}\}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5};w_{3})$. We will show that $\widehat{W}_{1}$ is an extended $5$-wheel. Note that $W_{1}$ and $W_{2}$ are full because they form a dwheel. This implies that $w_{3}$ is distinct from the vertices of $W_{2}$. It remains to show that $w_{3}$ does not form an edge with any of $w_{5}$, $v_{4}$, $v_{5}$ or $v_{1}$. That $w_{3}\nsim w_{5}$ follows by fullness of $W_{2}$. The remaining cases would each introduce a $4$-cycle:
  • $w_{3}\sim v_{4}$ introduces the $4$-cycle $(w_{3},v_{2},w_{5},v_{4})$.
  • $w_{3}\sim v_{5}$ introduces the $4$-cycle $(w_{3},v_{3},w_{5},v_{5})$.
  • $w_{3}\sim v_{1}$ introduces the $4$-cycle $(w_{3},v_{3},w_{5},v_{1})$.
The $5$-largeness of $X$ applied to any of these $4$-cycles would contradict fullness of $W_{1}$ or $W_{2}$. Thus $\widehat{W}_{1}$ is an extended $5$-wheel and so, by the $\widehat{W}_{5}$-condition, there exists a vertex $p$ of $X$ such that $\widehat{W}_{1}\subset X_{p}$. Fullness of $W_{2}$ implies that $p$ is distinct $w_{1}$ and $w_{2}$ so $p$ is distinct from the vertices of $W=W_{1}\cup W_{2}$. By a similar argument, we have that $\widehat{W}_{2}=W_{2}\cup\{v_{4}\}=(v_{2};w_{1},w_{2},w_{3},w_{4},w_{5};v_{4})$ is an extended $5$-wheel contained in $X_{q}$ for some vertex $q$ of $X\setminus W$. If $p=q$ then $W\subset X_{p}$ and we are done so we may assume that $p$ and $q$ are distinct. Thus we have a $4$-cycle $(p,q,w_{1},v_{1})$ to which we can apply $5$-largeness. Without loss of generality, we have $p\sim w_{1}$ so that we have a $4$-cycle $(p,w_{1},w_{2},w_{3})$. But $w_{1}\nsim w_{3}$ by fullness of $W_{2}$ so, applying $5$-largeness, we have $p\sim w_{2}$ and so $W=W_{1}\cup W_{2}\subset B_{1}(p)$. ∎
Claim 3.4.
If $W=W_{1}\cup W_{2}$ is a planar $(5,6)$-dwheel then $W$ is contained in a $1$-ball of $X$.
Proof.
Let $W_{1}\cup W_{2}=(w_{6};v_{1},v_{2},v_{3},v_{4},v_{5})\cup(v_{2};w_{1}=v_{1},w_{2},w_{3},$ $w_{4},w_{5}=v_{3},w_{6})$. Then $W_{1}$ and $W_{2}$ are full. Let $\widehat{W}_{1}=W_{1}\cup\{w_{2}\}=(w_{6};v_{1},v_{2},v_{3},v_{4},v_{5};w_{2})$. Let $\widehat{W}_{1}^{\prime}=W_{1}\cup\{w_{4}\}=(w_{6};v_{1},v_{2},v_{3},v_{4},v_{5};w_{4})$. See Figure 5.
$w_{6}$$v_{1}=w_{1}$$v_{2}$$v_{3}=w_{5}$$v_{4}$$v_{5}$$w_{2}$$w_{3}$$w_{4}$
Figure 5. A planar $(5,6)$-dwheel from the proof of Claim 3.4. The encircled vertices form an extended $5$-wheel $\widehat{W}_{1}$ contained in the link $X_{p}$ of the vertex $p$. The ensquared vertices form an extended $5$-wheel $\widehat{W}_{1}^{\prime}$ contained in $X_{q}$.
By fullness of $W_{2}$, we have $w_{2}\neq w_{6}$, $w_{2}\nsim w_{6}$ and $w_{2}\nsim v_{3}$. Then, to show that $\widehat{W}_{1}$ is an extended $5$-wheel, it suffices to show $w_{2}\nsim v_{4}$ and $w_{2}\nsim v_{5}$. But $w_{2}\sim v_{4}$ would introduce the $4$-cycle $(w_{2},v_{2},w_{6},v_{4})$ and $w_{2}\sim v_{5}$ would introduce the $4$-cycle $(w_{2},v_{2},w_{6},v_{5})$. Applying $5$-largeness to either of these $4$-cycles would contradict fullness of either $W_{1}$ or $W_{2}$. Thus $\widehat{W}_{1}$ is an extended $5$-wheel and so is contained in the link $X_{p}$ of a vertex $p$ of $X$. Fullness of $W_{2}$ ensures that $p$ is distinct from $w_{3}$ and $w_{4}$ so that $p\notin W=W_{1}\cup W_{2}$. By symmetry, the same argument can be applied to $\widehat{W}_{1}^{\prime}$ to show that $\widehat{W}_{1}^{\prime}\subset X_{q}$ for some vertex $q$ of $X\setminus W$.
If $p=q$ then we could apply $5$-largeness to the $4$-cycle $(w_{2},w_{3},w_{4},p)$ and then the fullness of $W_{2}$ would imply that $p\sim w_{3}$. Then we would have $W\subset X_{p}$, which would prove the claim. So we may assume that $p\neq q$. Then fullness of $W_{1}$ and $5$-largeness applied to the $4$-cycle $(v_{1},p,v_{3},q)$ implies that $p\sim q$.
$w_{6}$$v_{1}=w_{1}$$v_{2}$$v_{3}=w_{5}$$v_{4}$$v_{5}$$w_{2}$$w_{3}$$w_{4}$$p$$q$
Figure 6. The extended $5$-wheel $\widehat{W}$ from the proof of Claim 3.4.
Next we need to show that $\widehat{W}=(v_{2};p,w_{2},w_{3},w_{4},q;v_{5})$ is an extended $5$-wheel. See Figure 6. Fullness of $W_{1}$ implies that $v_{5}\neq v_{2}$ and $v_{5}\nsim v_{2}$. We need to show that $v_{5}$ is not adjacent to $w_{2}$, $w_{3}$ or $w_{4}$. But adjacency of $v_{5}$ to any of these vertices would introduce a $4$-cycle in $W$ two which applying $5$-largeness would either contradict fullness of $W_{1}$ or fullness of $W_{2}$. Thus $\widehat{W}$ is an extended $5$-wheel and so is contained in the link $X_{r}$ of a vertex $r$ of $X$. Fullness of $W_{1}$ and $W_{2}$ ensure that $r\notin W$.
Since $v_{5}\nsim v_{2}$, the $5$-largeness of $X$ applied to $4$-cycles $(v_{5},r,v_{2},v_{1})$ and $(v_{5},r,v_{2},w_{6})$ imply that $r$ is adjacent to $v_{1}$ and $w_{6}$. Then, similarly, the $4$-cycle $(w_{6},r,w_{4},v_{3})$ implies that $r\sim v_{3}$. Finally, the $4$-cycle $(v_{5},r,v_{3},v_{4})$ implies that $r\sim v_{4}$. Thus $W\subset B_{1}(r)$. ∎
The above three claims together imply that $X$ is $7$-located. ∎
4. A $\operatorname{CAT}(0)$ metric for $7$-located disc diagrams
In this section we show that the triangles of any $7$-located disc $D$ can be metrized in a natural way as Euclidean triangles such that $D$ is $\operatorname{CAT}(0)$. Only three isometry types of triangles are required and the choice of isometry type for each $2$-simplex $\sigma$ will depend only on the local combinatorics around $\sigma$.
Lemma 4.1.
Let $D$ be a $7$-located disc. Then any pair of adjacent internal vertices $u,v$ of $D$ satisfy $\deg(v)+\deg(w)\geq 12$.
Proof.
Since $D$ is a simplicial complex and $v$ is an interior vertex, it is the center of a full wheel $W_{v}$. Similarly, we have a full wheel $W_{w}$ in $D$ with center $w$ and $W_{v}\cup W_{w}$ is a planar dwheel of boundary length $\deg(v)+\deg(w)-4$. Since $D$ is planar, the dwheel $W_{v}\cup W_{w}$ cannot be contained in a $1$-ball. Thus $\deg(v)+\deg(w)-4\geq 8$. ∎
Definition 4.2.
Let $W$ be a $k$-wheel with central vertex $v$. We metrize each triangle of $W$ as a Euclidean triangle whose angle at $v$ is $\frac{2\pi}{k}$, whose remaining two angles are both equal to $\frac{(k-2)\pi}{2k}$ and whose boundary edge $e\subset\partial W$ has length $1$. We then metrize $W$ as a Euclidean polygonal complex. We call the resulting metric the flattened wheel metric on $W$.
Remark 4.3.
A $k$-wheel $W$ with the flattened wheel metric is isometric to a regular Euclidean $k$-gon of side length $1$. Under this isometry, the central vertex is sent to the center of the $k$-gon.
Definition 4.4.
Let $D$ be a simplicial disc such that for any $k_{1}$-wheel $W_{1}\subset D$ and any $k_{2}$-wheel $W_{2}\subset D$, if $k_{1}<6$ and $k_{2}<6$ then either $W_{1}=W_{2}$ or $W_{1}\cap W_{2}\subset\partial W_{1}\cap\partial W_{2}$. The flattened wheel metric on $D$ is the Euclidean polygonal complex metric obtained by metrizing each $k$-wheel of $D$ for which $k<6$ with the flattened wheel metric and metrizing any remaining triangles as regular Euclidean triangles of side length $1$. With this metric, the flattened wheels of $D$ are its $k$-wheels with $k<6$.
Lemma 4.5.
Let $D$ be a $7$-located disc. Let $W_{1}$ and $W_{2}$ be distinct wheels of $D$ of boundary length less than $6$. Then $W_{1}\cap W_{2}\subset\partial W_{1}\cap\partial W_{2}$. In particular, the flattened wheel metric is well-defined for $D$.
Proof.
If $W_{1}$ and $W_{2}$ intersect in their interiors then they must share a triangle $\sigma$. Thus the central vertices $v_{1}$ and $v_{2}$ of $W_{1}$ and $W_{2}$ are adjacent. But then, by Lemma 4.1, we have $|\partial W_{1}|+|\partial W_{2}|=\deg(v_{1})+\deg(v_{2})\geq 12$ so that one of the wheels must have boundary length at least $6$. ∎
In the next theorem we establish a connection between some $7$-located discs and $\operatorname{CAT}(0)$ spaces.
Theorem 4.6.
Let $D$ be a $7$-located simplicial disc endowed with the flattened wheel metric. Then $D$ is $\operatorname{CAT}(0)$.
Proof.
Since $D$ is simply connected, it suffices to show that the sum of the angles of the corners of triangles incident to any interior vertex $v$ of $D$ is at least $2\pi$ [3, Theorem II.5.4 and Lemma II.5.6]. Because $D$ is flag, it has no interior vertices of degree $3$. Any interior vertex of degree $4$ or $5$ is the central vertex of a flattened wheel and so has angle sum exactly $2\pi$. Away from flattened wheel centers, corner angles of triangles are either $\frac{\pi}{4}$, $\frac{3\pi}{10}$ or $\frac{\pi}{3}$. Thus any interior vertex of degree at least $8$ has angle sum at least $2\pi$.
Thus we may assume that $v$ is an interior vertex of degree $6$ or $7$. If $v$ is not incident to any flattened wheel then its incident triangles are all regular Euclidean and so have corner angle $\frac{\pi}{3}$. Thus $v$ has angle sum $\frac{6\pi}{3}=2\pi$ or $\frac{7\pi}{3}>2\pi$. So we may assume that $v$ is incident to a flattened wheel. Since $\deg(v)\leq 7$, by Lemma 4.1, any central vertex of a flattened wheel $W$ incident to $v$ has degree $5$. Then $v$ is incident to a vertex of degree $5$ and any triangle corner incident to $v$ has angle at least $\frac{3\pi}{10}$. So, by Lemma 4.1, we have $\deg(v)=7$ and so $v$ has angle sum at least $\frac{21\pi}{10}>2\pi$. ∎
Corollary 4.7.
There is a uniform quadratic function $f\colon\mathbb{N}\to\mathbb{N}$ such that any $7$-located disc $D$ of boundary length $|\partial D|=n$ has at most $f(n)$ triangles.
Proof.
The area of a minimal disc diagram for a loop of length $n$ in a $\operatorname{CAT}(0)$ space is bounded by the area of a circle of circumference $n$ in the Euclidean plane [3, Theorem III.2.17]. It follows that if $D$ is a $7$-located disc of boundary length $n$, it has area at most $\frac{n^{2}}{4\pi}$ when endowed with the flattened wheel metric. Each triangle of $D$ has area at least $\frac{1}{4}$ so the number of triangles of $D$ is at most $f(n)=\frac{n^{2}}{\pi}$. ∎
5. A Minimal Disc Diagram Lemma for $7$-located locally $5$-large complexes
In this section we prove that a minimal area disc diagram in a $7$-located locally $5$-large complex is $7$-located.
Remark 5.1.
The simplicial complex $X$ in Figure 7 is $7$-located but it is not locally $5$-large. However, it has a minimal area disc diagram $(D,f)$ that is not $7$-located. This example justifies the additional hypothesis of local $5$-largeness for the minimal disc diagram lemma.
Figure 7.
We shall refer to the following lemmas frequently in the rest of this section.
Lemma 5.2.
Let $X$ be a flag simplicial complex and let $\gamma$ be a homotopically trivial loop in $X$. Let $(D,f)$ be a disc diagram for $\gamma$. Let $(u,a,v,b)$ be a $4$-cycle in $X$. If $f(u)=f(v)$ then there exists a disc diagram $(D^{\prime},f^{\prime})$ for $\gamma$ of lesser area than $D$.
Proof.
The subdisc bounded by $(u,a,v,b)$ has at least two triangles. We delete this subdisc. We glue $u$ to $v$, the edge $\langle a,u\rangle$ to the edge $\langle a,v\rangle$, and the edge $\langle b,u\rangle$ to the edge $\langle b,v\rangle$. Thus we get a new disc $D^{\prime}$. See Figure 8. Let $f^{\prime}$ be the map induced by the gluing. This is well defined since $f(u)=f(v)$, $f(\langle a,u\rangle)=f(\langle a,v\rangle)$ and $f(\langle b,u\rangle)=f(\langle b,v\rangle)$. Thus $(D^{\prime},f^{\prime})$ is a disc diagram for $\gamma$ of lesser area than $D$.
Figure 8.
Lemma 5.3.
Let $X$ be a flag simplicial complex and let $\gamma$ be a homotopically trivial loop in $X$. Let $(D,f)$ be a disc diagram for $\gamma$. Let $u,v$ be vertices of a $5$-cycle $\alpha=(a,v,b,c,u)$ of $D$ with $f(u)=f(v)$. Then there exists a disc diagram $(D^{\prime},f^{\prime})$ for $\gamma$ of lesser area than $D$.
Proof.
The subdisc bounded by $\alpha$ has at least three $2$-simplices. We delete this subdisc. We obtain a new disc diagram $D^{\prime}$ by glueing $u$ to $v$, gluing the edge $\langle a,u\rangle$ to the edge $\langle a,v\rangle$ and gluing a $2$-simplex $\langle u,b,c\rangle$. See Figure 9. Let $f^{\prime}$ be the map induced by the gluing. This is well defined since $f(u)=f(v)$, $f(\langle a,u\rangle)=f(\langle a,v\rangle)$ and, by flagness, $\langle f(u),f(b),f(c)\rangle$ is a triangle in $X$. Thus $(D^{\prime},f^{\prime})$ is a disc diagram for $\gamma$ of lesser area than $D$.
Figure 9.
Lemma 5.4.
Let $X$ be a flag simplicial complex and let $\gamma$ be a homotopically trivial loop in $X$. Let $(D,f)$ be a disc diagram for $\gamma$. Let $u,v$ be vertices of a $6$-cycle $\alpha=(u,a,b,v,c,d)$ of $D$ with $d(u,v)=3$ such that $f(u)=f(v)$. Then there exists a disc diagram $(D^{\prime},f^{\prime})$ for $\gamma$ of lesser area than $D$.
Proof.
The subdisc bounded by $\alpha$ has at least four $2$-simplices. We obtain a new disc diagram $D^{\prime}$ by deleting the subdisc bounded by $\alpha$, gluing $u$ to $v$ and then gluing in two triangles $\langle a,b,u\rangle$ and $\langle u,c,d\rangle$. See Figure 10. Let $f^{\prime}$ be the map induced by gluing. This is well defined since $f(u)=f(v)$ and, by flagness, $\langle f(a),f(b),f(u)\rangle$, $\langle f(u),f(c),f(d)\rangle$ are triangles in $X$. Hence $(D^{\prime},f^{\prime})$ is a disc diagram for $\gamma$ of lesser area than $D$.
Figure 10.
Lemma 5.5.
Let $X$ be a flag and locally $5$-large simplicial complex. Let $(D,f)$ be a minimal area disc diagram for a cycle $\gamma$ in $X$. Then $D$ does not contain a $4$-wheel.
Proof.
Suppose $W$ is a $4$-wheel in $D$. If the restriction $f|_{W}$ is not injective then some pair of antipodal vertices $u,v\in\partial W$ have a common image $f(u)=f(v)$. This contradicts minimality by Lemma 5.2.
If the restriction $f|_{W}$ is injective then, since $X$ is locally $5$-large, a pair of antipodal vertices $u,v\in\partial W$ have images $f(u)$ and $f(v)$ that are joined by an edge $e$ in $X$. Since $X$ is flag, the image $f(\partial W)$ along with $e$ span two triangles so that $f(\partial W)$ has a disc diagram of area $2$. Thus we may remove the interior of $W$ in $D$ and replace it with an edge joining $u$ and $v$ and a pair of triangles to obtain a disc diagram for $\gamma$ of lesser area, contradicting minimality of $D$. ∎
Lemma 5.6.
Let $X$ be a flag and locally $5$-large simplicial complex. Let $(D,f)$ be a minimal area disc diagram for a cycle $\gamma$ in $X$ and let $W$ be a $5$-wheel in $D$. Then the image $f(W)$ is a full $5$-wheel of $X$.
Proof.
It follows from Lemma 5.3 that $f|_{W}$ is injective. Then $f(W)$ is a $5$-wheel in $X$. We need only verify that it is full. If not then some pair of vertices $u,v\in\partial W$ that are not adjacent in $\partial W$ have images $f(u)$ and $f(v)$ that are joined by an edge $e$. Since $X$ is flag, the edge $e$ spans a triangle with each of $f(u)$ and $f(v)$. Let $a$ be the unique vertex of $\partial W$ adjacent to both $u$ and $v$. Let $c$ be the central vertex of $W$. Then $(u,a,v,c)$ is a $4$-cycle bounding a subdisc of $D$ with two triangles. We delete this subdisc, join $u$ and $v$ by an edge $e^{\prime}$ and add triangles $\langle a,u,v\rangle$ and $\langle c,u,v\rangle$ to obtain a new disc diagram $D^{\prime}$ for $\gamma$ of the same (minimal) area. In particular, $D^{\prime}$ is a minimal area disc diagram. But the triangle $\langle c,u,v\rangle$ along with the three remaining original triangles of $W$ form a $4$-wheel, which contradicts minimality by Lemma 5.5. ∎
Lemma 5.7.
Let $X$ be a $7$-located locally $5$-large simplicial complex. Let $(D,f)$ be a minimal area disc diagram for a cycle $\gamma$ in X. Let $W$ be a $(5,5)$- or $(5,6)$-dwheel of $D$. Then the restriction $f|_{W}$ is not injective.
Proof.
For the sake of deriving a contradiction, we assume that $f|_{W}$ is injective. Since $D$ is planar, the dwheel $W$ is also planar. We consider first the case where $W$ is a planar $(5,5)$-dwheel. By Lemma 5.6, the $(5,5)$-dwheel $f(W)$ has full wheels. Then, since $X$ is $7$-located, the image $f(W)$ is contained in the $1$-ball centered at a vertex $v$ of $X$. Since $X$ is flag, it follows that $f(\partial W)$ has a disc diagram with at most $6$ triangles, contradicting minimality of $D$.
$w_{6}$$v_{1}=w_{1}$$v_{2}$$v_{3}=w_{5}$$v_{4}$$v_{5}$$w_{2}$$w_{3}$$w_{4}$
Figure 11. A planar $(5,6)$-dwheel.
We now consider the case where $W$ is a planar $(5,6)$-dwheel. Let $W=W_{1}\cup W_{2}=(w_{6};v_{1},v_{2},v_{3},v_{4},v_{5})\cup(v_{2};w_{1}=v_{1},w_{2},w_{3},$ $w_{4},w_{5}=v_{3},w_{6})$; see Figure 11. By Lemma 5.6, the image $f(W_{1})$ is a full $5$-wheel. If the $6$-wheel $f(W_{2})$ is also a full then we can argue as in the $(5,5)$-dwheel case that $f(\partial W)$ has a disc diagram with at most $7$ triangles, which leads to a contradiction with the minimality of $D$. Thus we need only consider the case where $f(W_{2})$ is not full. Then some pair of vertices $u,v\in\partial W_{2}$ that are not adjacent in $\partial W_{2}$ have images $f(u)$ and $f(v)$ that are adjacent.
Claim 5.8.
Either $\{u,v\}=\{w_{6},w_{2}\}$ or $\{u,v\}=\{w_{6},w_{4}\}$.
Proof.
We first rule out the possibility that $u$ and $v$ are antipodal in $\partial W_{2}$. If this were so, we could cut $D$ open along the path $(u,v_{2},w)$ and fill the resulting boundary path with an edge joining $u$ and $w$ (mapping to the edge $\langle f(u),f(v)\rangle$) and a pair of triangles (which map to triangles in $X$ by flagness). In the resulting disc diagram for $\gamma$, the cycle $\partial W_{2}$ bounds a disc with two $4$-wheels. But, by arguments similar to those in the proof of Lemma 5.5, each $4$-wheel boundary (as a cycle in $X$) has a disc diagram with at most two triangles. Thus $\partial W_{2}$ has a disc diagram with at most four triangles, contradicting minimality of $(D,f)$.
Thus $u$ and $v$ are at distance $2$ in $\partial W_{2}$. Since $f(W_{1})$ is full in $X$, we have $\{u,v\}\neq\{w_{1},w_{5}\}$. It remains to rule out the remaining possibilities: $\{u,v\}=\{w_{1},w_{3}\}$, $\{u,v\}=\{w_{2},w_{4}\}$ and $\{u,v\}=\{w_{3},w_{5}\}$. In these cases, we perform a disc diagram surgery which turns $v_{2}$ into the center of a $5$-wheel, similar to the surgery performed in the proof of Lemma 5.6 that produced a $4$-wheel. The resulting disc diagram has the same (minimal) area but has a $5$-wheel and thus the argument reduces to the $(5,5)$-dwheel case. ∎
Thus the additional edges in the full subcomplex induced on the $6$-wheel $f(W_{2})$ are either $\langle f(w_{6}),f(w_{2})\rangle$ or $\langle f(w_{6}),f(w_{4})\rangle$ or both. If just one of these is present then $f(W)$ spans a non-planar $(5,5)$-dwheel with full wheels so that $\partial W$ has a disc diagram with at most $7$ triangles, contradicting minimality of $(D,f)$. If, on the other hand, both edges are present then we perform two disc diagram surgeries as in the proof of Lemma 5.5 transforming $W$ from a $(5,6)$-dwheel to a $(7,4)$-dwheel. The resulting disc diagram has the same (minimal) area and yet contains a $4$-wheel, contradicting Lemma 5.5. ∎
We prove next the minimal disc diagram lemma for $7$-located locally $5$-large simplicial complexes.
Theorem 5.9.
Let $X$ be a $7$-located locally $5$-large simplicial complex. Let $\gamma$ be a homotopically trivial loop in $X$. Any minimal area disc diagram $(D,f)$ for $\gamma$ is $7$-located and locally $5$-large.
Proof.
By Lemma 5.5, there are no $4$-wheels in $D$. It follows that $D$ is locally $5$-large so we need only prove that $D$ is $7$-located. Suppose for the sake of finding a contradiction that $D$ is not $7$-located. Then $D$ has a dwheel $W=W_{1}\cup W_{2}$ of boundary length at most 7 that is not contained in a $1$-ball. Since $D$ is planar, so is $W$. Since $D$ has no $4$-wheels, the dwheel $W$ must be a planar $(5,5)$- or $(5,6)$-dwheel. By Lemma 5.7, the restriction $f|_{W}$ is not injective. Then there exist distinct vertices $v,w\in W$ such that $f(v)=f(w)$.
We consider first the case where $v$ and $w$ belong to a common wheel $W_{i}$ of $W$. Since $f$ is a nondegenerate map and $X$ is simplicial, the vertices $v$ and $w$ cannot be adjacent. Thus they must be contained in the boundary $\partial W_{i}$ and be at distance $2$ or $3$ in $\partial W_{i}$. Note that $W_{i}$ is necessarily a $6$-wheel in the case that $d_{\partial W_{i}}(v,w)=3$. But then we contradict minimality by Lemma 5.2 if $d_{\partial W_{i}}(v,w)=2$ or by Lemma 5.4 if $d_{\partial W_{i}}(v,w)=3$.
We consider now the case where $v$ and $w$ do not belong to a common wheel of $W$. It will be helpful here to name the vertices of $W$ so let $W=W_{1}\cup W_{2}=(w_{k};v_{1},v_{2},\dots,v_{5})\cup(v_{2};w_{1}=v_{1},w_{2},\dots,w_{k-1}=v_{3},w_{k})$ with $k\in\{5,6\}$. We first consider the $(5,5)$-dwheel case, i.e., the case where $k=5$. See Figure 3. Up to symmetry, there are only two cases to consider: $(v,w)=(v_{4},w_{3})$ and $(v,w)=(v_{4},w_{2})$. In the case $(v,w)=(v_{4},w_{3})$, the $5$-cycle $(v_{4},v_{3},w_{3},v_{2},w_{5})$ contradicts minimality by Lemma 5.3. In the case $(v,w)=(v_{4},w_{2})$, the $6$-cycle $\partial W=(v_{4},v_{3},w_{3},w_{2},w_{1},v_{5})$ contradicts minimality by Lemma 5.4.
We now consider the $(5,6)$-dwheel case, i.e., the case where $k=6$. See Figure 11. Up to symmetry, there are three cases to consider: $(v,w)=(v_{4},w_{4})$, $(v,w)=(v_{4},w_{3})$ and $(v,w)=(v_{4},w_{2})$. As with the $(5,5)$-dwheel, in each case we contradict minimality by applying either Lemma 5.3 to a $5$-cycle or Lemma 5.4 to a $6$-cycle. Specifically, for the case $(v,w)=(v_{4},w_{4})$ we consider the $5$-cycle $(v_{4},v_{3},w_{4},v_{2},w_{6})$, for the case $(v,w)=(v_{4},w_{3})$ we consider the $6$-cycle $(v_{4},v_{3},w_{4},w_{3},v_{2},w_{6})$ and for the $(v,w)=(v_{4},w_{2})$ case we consider the $6$-cycle $(v_{4},w_{6},v_{2},w_{2},w_{1},v_{5})$.